# Compounding Vulnerability: Hub Removal Can Trigger Cascade Phase Transitions While Degrading Percolation Robustness in Barabási–Albert Networks

### 07 Mar 2026 v2

@@ I Introduction
Prior work has shown qualitatively that complex networks can be “more robust and more fragile” depending on the stress model [10, 15], and recent studies have explored cascading failures across interdependent networks [28, 35] and scaling laws of failure dynamics [29], but stopped short of characterizing the mechanism or providing a predictive condition for when compounding vulnerability occurs under deliberate intervention on a single network. Here we address this gap directly, studying the canonical intervention of hub removal (top 10% of nodes by degree) in Barabási–Albert networks [9] and measuring its joint effect on percolation and cascade dynamics.
Our contributions are threefold:
@@
    A systematic study of the joint percolation-cascade phase structure under hub removal in BA networks, demonstrating that both stress models can worsen simultaneously at intermediate cascade thresholds—a phenomenon that neither stress model alone would predict.
@@
    A controlled experimental decomposition showing that hub-mediated cascade suppression is primarily dynamical (threshold-based) rather than structural (connectivity-based), demonstrated by the ${\sim}95\%/{\sim}19\%/{\sim}0.3\%$ cascade split across hub vulnerability conditions on the same network topology.
@@
    A closed-form analytical derivation of the post-removal cascade branching factor $z_{1}(\varphi,\alpha,m)$ under the configuration-model approximation, providing a predictive expression for cascade window expansion under targeted hub removal.
@@ II.1 Network Models
Barabási–Albert (BA): Generated by preferential attachment with $m=2$ new edges per node, yielding $P(k)\propto k^{-3}$. Degree heterogeneity $\kappa=\langle k^{2}\rangle/\langle k\rangle\approx 10$1313. We use N=500N=50010,00010{,}000 for finite-size scaling and N=2, [^0] N=2{,}000 for the $\varphi$-sweep.
@@ III.1 Experimental Design
Conditions B and B2 modify only the activation thresholds of hub nodes—the network topology is identical to baseline. This cleanly separates the dynamical firewall mechanism from connectivity effects.
@@ III.3 Decomposition of Effects
Pure vulnerability effect (B vs. A). Making hubs easy to activate—without changing the network topology—produces cascades of 95.0%, far exceeding the 18.7% seen under actual hub removal. When hubs become vulnerable, their high degree transforms them from cascade barriers into cascade amplifiers: a hub with $k=100$ that activates with a single neighbor immediately exposes 100 neighbors to cascade pressure. The cascade difference between A (0.28%) and B (95.0%) is entirely dynamical.
@@ IV.1 Three Regimes
Regime II ($0.20<\varphi\leq 0.25$): Phase transition zone. Hub removal opens a cascade window where the pre-removal network is subcritical but the post-removal network is supercritical. The boundary at $\varphi\approx 0.20$ is set by the integer threshold $\lfloor 1/\varphi\rfloor$ crossing from 5 to 4 (Table 6). At $\varphi=0.22$, mean cascade size increases from $0.86\%$ [95% CI: 0.43–1.30] to $23.1\%$ [95% CI: 21.3–24.9] ($n_{\mathrm{eff}}=50$ independent networks, 20 seeds each; total n=1,000n=1{,}000; within-network seeds are correlated, so network-level effective sample size is $n_{\mathrm{eff}}\approx 50$). This constitutes a genuine phase transition, confirmed by bimodal cascade size distributions (Fig. 2): the probability of a global cascade ($C>10\%$) rises from 0/1,000 before removal to 377/1,000 after removal ($P_{\mathrm{global}}=0.377$, Wilson 95% CI: $[0.348,0.406]$). Post-removal trials split between a subcritical mode ($C<1\%$) and a supercritical mode ($C>10\%$), with few intermediate values—the hallmark of a system straddling the critical point. The hub vulnerability experiment (Table 2, $n=200$: 10 networks $\times$ 20 seeds) yields directionally consistent results ($0.28\%\to 18.7\%$); the lower baseline (0.28% vs. 0.86%) reflects higher variance at small effective sample ($n_{\mathrm{eff}}=10$ networks) and is order-of-magnitude consistent with the main experiment.
Figure 2: Cascade size distribution after hub removal at $\varphi=0.22$, BA(N=2, [^0] N=2{,}000, $m=2$, 200 trials). The main panel shows the post-removal distribution, which is bimodal with modes at $1\%$ (subcritical) and $10\%$ (global cascade). The inset shows the pre-removal distribution, which is unimodal near zero—all cascades subcritical. The transition from unimodal to bimodal is the signature of a phase transition straddling the critical point.
@@ 4 ${\sim}17$ 7/10 10/10 10/10 10/10
IV.6 Finite-Size Scaling
@@
The N=2, [^0] N=2{,}000 estimate ($\Delta F_{p}=+352\%$) reported throughout is therefore conservative: larger networks show even stronger percolation degradation, consistent with finite-size corrections decaying as $O(1/\ln N)$. The narrowing standard deviations confirm convergence. The cascade phenomenon persists across N=500N=50010,00010{,}000, with the transition sharpening as $N$ increases (see repository datasets and scripts in Appendix E).
@@ V.3 Derivation of $z_{1}$ After Hub Removal
where $K^{*}=\min(K_{\text{eff}},k_{\text{cut}})$ and $H_{n}$ is the $n$th harmonic number. This approximation systematically overestimates $z_{1}$ by 17–36% because it uses the original branching structure k(k1)k(k-1) rather than the thinned post-removal degrees; it should be understood as an upper bound on the true post-removal $z_{1}$. The exact binomial computation in Table 6 does not use this approximation and provides the accurate values reported throughout.
@@ V.4 Numerical Verification of $z_{1}$
Sensitivity to $\rho$ estimation. The configuration-model approximation gives $\rho=0.375$, while direct measurement on BA networks yields $\rho_{\mathrm{emp}}=0.421\pm 0.011$ (12.3% higher), reflecting degree–degree correlations absent in the configuration model. Using $\rho_{\mathrm{emp}}$, the post-removal $z_{1}$ at $\varphi=0.22$ decreases from 1.195 to 1.140—still supercritical ($z_{1}>1$). The crossing $z_{1}^{\mathrm{pre}}<1<z_{1}^{\mathrm{post}}$ is preserved, confirming that the cascade phase transition prediction is robust to the $\rho$ underestimation inherent in the configuration-model approximation.
@@ VI.1 The Dual-Function Constraint
Our results reveal that within the intervention class of uniform node deletion on static, unweighted networks, hub nodes simultaneously provide percolation robustness (via their $k^{2}$ contribution to $\langle k^{2}\rangle$, keeping $p_{c}$ low) and cascade resistance (via their high activation barrier $k\cdot\varphi\gg 1$, blocking cascade propagation). Both functions are monotonically linked through the single structural parameter $k$. Across all tested networks and parameter combinations ($\varphi\in[0.15,0.28]$, $\alpha=0.10$, BA/ER/WS topologies, and four real-world networks), we observe no Pareto improvement: uniform hub deletion worsens both metrics simultaneously in every case examined. We note that this is an empirical observation, not a proven impossibility result; whether alternative intervention classes—degree-preserving rewiring, weighted edge adjustment, or adaptive threshold modification—can decouple these functions is an important open question for future work.
@@ VI.2 Relation to Prior Work
Gleeson and Cahalane (2007) [12] derived the cascade branching condition $z_{1}>1$ for random networks. We extend this to the post-removal setting, deriving a closed-form $z_{1}$ under the configuration-model approximation for targeted hub removal (§V.3) and demonstrating that the $z_{1}$ crossing explains the observed phase transition. Importantly, the cascade window expansion formula $\varphi^{*}_{\mathrm{post}}/\varphi^{*}_{\mathrm{pre}}\approx 1/(1-\rho)$ is not specific to the BA model: for any degree distribution, $\rho$ measures the edge fraction incident to removed hubs, and the expansion factor $1/(1-\rho)$ follows from the configuration-model $z_{1}$ framework. In homogeneous networks (Poisson, exponential), $\rho\ll 1$ and the expansion is negligible; in heterogeneous networks (power-law, BA), $\rho\sim 0.3$$0.5$ and the expansion is substantial. The degree heterogeneity threshold $\kappa>{\sim}10$ is thus a proxy for $\rho$ being large enough to produce an observable phase transition, not a fundamental constant of the BA model.
Albert, Jeong and Barabási (2000) [10] established the “robust yet fragile” duality for scale-free networks under percolation. We extend this to encompass threshold cascade dynamics, showing that the duality becomes compounding vulnerability under hub removal—both fragilities compound rather than trade off.
@@ VI.3 Real-World Network Validation
To validate beyond synthetic BA networks, we test on four real-world networks from Stanford SNAP with diverse degree distributions and clustering coefficients, applying the same hub removal protocol ($\alpha=0.10$) and cascade sweep ($\varphi\in[0.05,0.40]$, 40 seeds per $\varphi$). We estimate $p_{c}$ using both the susceptibility-peak and $S=0.5$ methods.
Table 7: Real-world network validation of compounding vulnerability under hub removal ($\alpha=0.10$). $\Delta F_{p}$: percolation fragility change (susceptibility peak). $C_{\max}^{\mathrm{pre/post}}$: peak mean cascade size before/after hub removal across all $\varphi$ values tested.
@@
Bitcoin OTC has $\gamma<2$; the finite-$\kappa$ analytical framework is not strictly applicable. $S=0.5$ method used where susceptibility peak is unreliable. High clustering ($C=0.622$); susceptibility peak yields anomalous results.
@@
Compounding vulnerability (both percolation and cascade worsening) is cleanly confirmed in p2p-Gnutella ($\kappa=17.7$, $C=0.011$; $\Delta F_{p}=+75\%$ by susceptibility peak) and Wiki-Vote ($\kappa=145.4$, $C=0.142$; $\Delta F_{p}=+594\%$). For Email-Enron ($C=0.497$) and CA-HepPh ($C=0.622$), the susceptibility-peak method gives anomalous percolation results ($\Delta F_{p}=-6.8\%$ and $80.4\%$ respectively), while the $S=0.5$ method yields $\Delta F_{p}=+31\%$ and $58\%$. We report $S=0.5$ values for these two networks (marked $\ddagger$} in Table 7). This estimator disagreement arises in high-clustering networks where the susceptibility landscape has multiple peaks; it does not affect the cascade amplification, which is robust across all estimators and all four networks. However, we note that full compounding vulnerability is confirmed with a single estimator only for the two low-clustering networks.
@@ VI.5 Limitations
(i) We validate on synthetic networks with controlled degree heterogeneity and on multiple real-world networks from Stanford SNAP. The Bitcoin OTC Trust Network (N=5,881N=5{,}881, $\gamma\approx 1.9$) confirms the directional effect but, with $\gamma<2$ implying divergent $\langle k^{2}\rangle$, the finite-$\kappa$ analytical framework of §V is not strictly applicable to this network; the agreement is empirical rather than analytical. (ii) The bond percolation threshold $p_{c}$ is estimated using the susceptibility peak method (maximum of the mean finite-cluster size), the standard approach in computational percolation [7]. We verify robustness by comparing against the $S=0.5$ crossing and the Molloy–Reed analytical estimate; all three methods confirm $\Delta F_{p}\gg 0$ (Appendix A). (iii) We test sensitivity to threshold heterogeneity using Beta-distributed per-node thresholds centered on $\varphi$ (Appendix D); the compounding vulnerability effect persists under moderate and high heterogeneity, with amplification ranging from $349\%$ to $1{,}955\%$ at $\varphi=0.22$. (iv) The configuration-model approximation in §V.3 ignores BA disassortativity and the correlation structure induced by targeted removal. Appendix C validates the $z_{1}$ framework on configuration-model networks (where it is exact by construction), confirming that the approximation captures the dominant mechanism and that BA-specific correlations partially dampen rather than create the effect. (v) Static networks are studied; adaptive rewiring could mitigate or amplify the effect. (vi) The Watts cascade model is one of several cascade mechanisms; generalization to load-redistribution [18], cascade-based attacks [21], and DebtRank [14] models is an important direction. (vii) The $\kappa>{\sim}10$ threshold is derived for BA networks; the precise threshold depends on both the degree distribution shape and the cascade threshold $\varphi$, and should be understood as a BA-specific onset condition rather than a distribution-independent constant.

### 12 Mar 2026 v3

# Compounding Vulnerability: Simultaneous Degradation of Percolation and Cascade Robustness Under Targeted Hub Removal

@@ I Introduction
Prior work has shown qualitatively that complex networks can be “more robust and more fragile” depending on the stress model [10, 15], and recent studies have explored cascading failures across interdependent networks [28, 35] and scaling laws of failure dynamics [29]. The qualitative direction of the joint effect—that hub removal should worsen both percolation (via $k^{2}$ loss) and cascades (via stable-node removal)—follows from these classical results. What has been missing is a quantitative characterization: where in parameter space does compounding vulnerability occur, how large is the effect, and what mechanism dominates? Here we provide this characterization for the canonical intervention of hub removal (top 10% of nodes by degree) in Barabási–Albert networks [9], measuring its joint effect on percolation and cascade dynamics.
Our contributions are:
@@
    A quantitative mapping of the joint percolation-cascade phase structure under hub removal in BA networks, identifying the parameter window $\varphi\in(0.20,0.26)$ where hub removal shifts the cascade dynamics from subcritical to supercritical while simultaneously degrading percolation.
@@
    A controlled experimental decomposition—including a neutral-threshold condition ($\varphi_{\mathrm{hub}}=1/k_{i}$) that isolates firewall removal from amplifier effects—showing that hub-mediated cascade suppression is primarily dynamical (threshold-based) rather than structural (connectivity-based).
@@
    A closed-form expression for the post-removal cascade branching factor $z_{1}(\varphi,\alpha,m)$ under the configuration-model approximation, providing a predictive condition for cascade window expansion that is robust to the 12% $\rho$-estimation error introduced by BA degree correlations.
@@ II.1 Network Models
Barabási–Albert (BA): Generated by preferential attachment with $m=2$ new edges per node, yielding $P(k)\propto k^{-3}$. Degree heterogeneity $\kappa=\langle k^{2}\rangle/\langle k\rangle\approx 10$1313. We use N=500N=50010,00010{,}000 for finite-size robustness checks and N=2, [^0] N=2{,}000 for the $\varphi$-sweep.
@@ III.1 Experimental Design
We additionally test Bn (neutral): hubs set to $\varphi_{\mathrm{hub},i}=1/k_{i}$ (single-neighbor activation without extra amplification), which isolates firewall removal from potential amplifier effects—the exact control proposed by the amplifier-confound concern described in Limitation (viii). Conditions B, Bn, and B2 modify only the activation thresholds of hub nodes—the network topology is identical to baseline. This cleanly separates the dynamical firewall mechanism from connectivity effects.
@@ III.2 Results
Table 2 summarizes the cascade outcomes under each condition.
@@ III.3 Decomposition of Effects
Pure vulnerability effect (B, Bn vs. A). Making hubs easy to activate—without changing the network topology—produces cascades of 95.0% (B) and 96.5% (Bn), far exceeding the 18.7% seen under actual hub removal. Crucially, conditions B and Bn yield statistically indistinguishable cascade sizes ($p>0.3$, two-sample $t$-test on network means), demonstrating that the cascade explosion is due to firewall removal, not super-amplification. For BA hubs with degrees $k\in[7,60+]$, both $\varphi_{\mathrm{hub}}=0.01$ and $\varphi_{\mathrm{hub}}=1/k_{i}$ yield the same discrete threshold $\lceil k\cdot\varphi_{\mathrm{hub}}\rceil=1$, explaining the equivalence. The cascade difference between A (0.28%) and Bn (96.5%) is entirely dynamical and free of the amplifier confound.
@@ IV.1 Three Regimes
Regime II ($0.20<\varphi\leq 0.25$): Phase transition zone. Hub removal opens a cascade window where the pre-removal network is subcritical but the post-removal network is supercritical. The boundary at $\varphi\approx 0.20$ is set by the integer threshold $\lfloor 1/\varphi\rfloor$ crossing from 5 to 4 (Table 6). At $\varphi=0.22$, mean cascade size increases from $0.86\%$ [95% CI: 0.43–1.30] to $23.1\%$ [95% CI: 21.3–24.9] (n=1,000n=1{,}000 total trials from 50 independent networks $\times$ 20 seeds each; since within-network seeds share topology, we treat each network mean as one independent observation, yielding a conservative effective sample size $n_{\mathrm{eff}}=50$). This constitutes a genuine phase transition, confirmed by bimodal cascade size distributions (Fig. 2): the probability of a global cascade ($C>10\%$) rises from 0/1,000 before removal to 377/1,000 after removal ($P_{\mathrm{global}}=0.377$, Wilson 95% CI: $[0.348,0.406]$). Post-removal trials split between a subcritical mode ($C<1\%$) and a supercritical mode ($C>10\%$), with few intermediate values—the hallmark of a system straddling the critical point. The hub vulnerability experiment (Table 2, $n=200$: 10 networks $\times$ 20 seeds) yields directionally consistent results ($0.28\%\to 18.7\%$); the lower baseline (0.28% vs. 0.86%) reflects higher variance at small effective sample ($n_{\mathrm{eff}}=10$ networks) and is order-of-magnitude consistent with the main experiment.
Figure 2: Cascade size distribution after hub removal at $\varphi=0.22$, BA(N=2, [^0] N=2{,}000, $m=2$). The histogram shows 200 independent trials (10 networks $\times$ 20 seeds); the upgraded experiment (n=1,000n=1{,}000; 50 networks $\times$ 20 seeds) yields a consistent global-cascade rate $P_{\mathrm{global}}=377/1{,}000=0.377$ (Wilson 95% CI: $[0.348,0.406]$). The main panel shows the post-removal distribution, which is bimodal with modes at $1\%$ (subcritical) and $10\%$ (global cascade). The inset shows the pre-removal distribution, which is unimodal near zero—all cascades subcritical. The transition from unimodal to bimodal is the signature of a phase transition straddling the critical point.
@@ 4 ${\sim}17$ 7/10 10/10 10/10 10/10
IV.6 Finite-Size Robustness
@@
The N=2, [^0] N=2{,}000 estimate ($\Delta F_{p}=+352\%$) reported throughout is therefore conservative: larger networks show even stronger percolation degradation, consistent with finite-size corrections decaying as $O(1/\ln N)$. We note that $\Delta F_{p}$ is $N$-dependent because $p_{c}^{\mathrm{pre}}$ decreases with $N$ (reflecting growing $\kappa$); absolute $p_{c}$ values in Table 5 provide the $N$-invariant characterization. The narrowing standard deviations confirm convergence. The cascade phenomenon persists across N=500N=50010,00010{,}000, with the transition sharpening as $N$ increases (see repository datasets and scripts in Appendix E).
@@ V.3 Derivation of $z_{1}$ After Hub Removal
where $K^{*}=\min(K_{\text{eff}},k_{\text{cut}})$ and $H_{n}$ is the $n$th harmonic number. This approximation systematically overestimates $z_{1}$ by 17–36% because it uses the original branching structure k(k1)k(k-1) rather than the thinned post-removal degrees; it should be understood as an upper bound on the true post-removal $z_{1}$. The full binomial computation in Table 6 does not use this mean-field approximation and provides the configuration-model predictions reported throughout; these remain approximations for BA networks due to degree correlations (see §V.4).
@@ V.4 Numerical Verification of $z_{1}$
Sensitivity to $\rho$ estimation. The configuration-model approximation gives $\rho=0.375$, while direct measurement on BA networks yields $\rho_{\mathrm{emp}}=0.421\pm 0.011$ (12.3% higher), reflecting degree–degree correlations absent in the configuration model. Using $\rho_{\mathrm{emp}}$, the post-removal $z_{1}$ at $\varphi=0.22$ decreases from 1.195 to 1.140—still supercritical ($z_{1}>1$). Crucially, the cascade window boundary is also robust: at $\varphi=0.25$ (the upper edge of Regime II), $z_{1}^{\mathrm{post}}(\rho_{\mathrm{emp}})=1.140>1$, and the window closes only at $\varphi\geq 0.26$ under both $\rho$ estimates. The crossing $z_{1}^{\mathrm{pre}}<1<z_{1}^{\mathrm{post}}$ is preserved across the entire Regime II parameter range, confirming that the cascade phase transition prediction is robust to the $\rho$ underestimation inherent in the configuration-model approximation.
@@ VI.1 The Dual-Function Constraint
Our results reveal that within the intervention class of uniform node deletion on static, unweighted networks, hub nodes simultaneously provide percolation robustness (via their $k^{2}$ contribution to $\langle k^{2}\rangle$, keeping $p_{c}$ low) and cascade resistance (via their high activation barrier $k\cdot\varphi\gg 1$, blocking cascade propagation). Both functions are monotonically linked through the single structural parameter $k$. Across all tested networks and parameter combinations ($\varphi\in[0.15,0.28]$, $\alpha=0.10$, BA/ER/WS topologies, and five real-world networks), we observe no Pareto improvement: uniform hub deletion worsens both metrics simultaneously in every case examined. We note that this is an empirical observation, not a proven impossibility result; whether alternative intervention classes—degree-preserving rewiring, weighted edge adjustment, or adaptive threshold modification—can decouple these functions is an important open question for future work.
We emphasize that the qualitative direction of compounding vulnerability—that hub removal should worsen both percolation and cascades—is expected from classical results: Watts (2002) identified high-degree nodes as cascade barriers, and Albert et al. (2000) showed that hub removal degrades connectivity. The contribution of the present work is the quantitative characterization: the precise location of the $\varphi$-window via the post-removal $z_{1}$, the controlled decomposition of dynamical vs. topological mechanisms (with the Bn condition confirming negligible amplifier effects), and the identification of the degree-heterogeneity threshold $\kappa>{\sim}10$ as the onset condition.
@@ VI.2 Relation to Prior Work
Gleeson and Cahalane (2007) [12] derived the cascade branching condition $z_{1}>1$ for random networks; Melnik et al. [42] and Hackett et al. [43] extended this to multi-edge effects and small-world cascades respectively. We extend the framework to the post-removal setting, deriving a closed-form $z_{1}$ under the configuration-model approximation for targeted hub removal (§V.3), using the binomial edge-thinning technique of Newman [44], and demonstrating that the $z_{1}$ crossing explains the observed phase transition. Importantly, the cascade window expansion formula $\varphi^{*}_{\mathrm{post}}/\varphi^{*}_{\mathrm{pre}}\approx 1/(1-\rho)$ is not specific to the BA model: for any degree distribution, $\rho$ measures the edge fraction incident to removed hubs, and the expansion factor $1/(1-\rho)$ follows from the configuration-model $z_{1}$ framework. In homogeneous networks (Poisson, exponential), $\rho\ll 1$ and the expansion is negligible; in heterogeneous networks (power-law, BA), $\rho\sim 0.3$$0.5$ and the expansion is substantial. The degree heterogeneity threshold $\kappa>{\sim}10$ is thus a proxy for $\rho$ being large enough to produce an observable phase transition, not a fundamental constant of the BA model.
Albert, Jeong and Barabási (2000) [10] established the “robust yet fragile” duality for scale-free networks under percolation, a phenomenon predicted more broadly by the Highly Optimized Tolerance (HOT) framework of Carlson and Doyle [41], who showed that systems optimized for one failure mode are generically fragile to others. We extend this duality to encompass threshold cascade dynamics, showing that it becomes compounding vulnerability under hub removal—both fragilities compound rather than trade off. Dorogovtsev et al. [40] provide a comprehensive review of critical phenomena on complex networks, within which our results identify a specific compounding mechanism at the intersection of percolation and cascade transitions.
@@ VI.3 Real-World Network Validation
To validate beyond synthetic BA networks, we test on five real-world networks from Stanford SNAP with diverse degree distributions and clustering coefficients, applying the same hub removal protocol ($\alpha=0.10$) and cascade sweep ($\varphi\in[0.05,0.40]$, 40 seeds per $\varphi$). We estimate $p_{c}$ using both the susceptibility-peak and $S=0.5$ methods.
Table 7: Real-world network validation of compounding vulnerability under hub removal ($\alpha=0.10$) across five SNAP networks. $\Delta F_{p}$: percolation fragility change (susceptibility peak unless noted). $C_{\max}^{\mathrm{pre/post}}$: peak mean cascade size before/after hub removal across all $\varphi$ values tested.
@@
Bitcoin OTC has $\gamma<2$; the finite-$\kappa$ analytical framework is not strictly applicable. $S=0.5$ method used where susceptibility peak is unreliable. High clustering; susceptibility peak yields anomalous results; $S=0.5$ method used.
@@
Compounding vulnerability (both percolation and cascade worsening) is cleanly confirmed in p2p-Gnutella ($\kappa=17.7$, $C=0.011$; $\Delta F_{p}=+75\%$ by susceptibility peak) and Wiki-Vote ($\kappa=145.4$, $C=0.142$; $\Delta F_{p}=+594\%$). For Email-Enron ($C=0.497$) and CA-HepPh ($C=0.622$), the susceptibility-peak method gives anomalous percolation results ($\Delta F_{p}=-6.8\%$ and $80.4\%$ respectively), while the $S=0.5$ method yields $\Delta F_{p}=+31\%$ and $58\%$. We report $S=0.5$ values for these two networks (marked $\ddagger$} in Table 7). This estimator disagreement arises in high-clustering networks where the susceptibility landscape has multiple peaks; it does not affect the cascade amplification, which is robust across all estimators and all five networks. However, we note that full compounding vulnerability is confirmed with a single estimator only for the two low-clustering networks.
@@ VI.5 Limitations
(i) We validate on synthetic networks with controlled degree heterogeneity and on multiple real-world networks from Stanford SNAP. The Bitcoin OTC Trust Network (N=5,881N=5{,}881, $\gamma\approx 1.9$) confirms the directional effect but, with $\gamma<2$ implying divergent $\langle k^{2}\rangle$, the finite-$\kappa$ analytical framework of §V is not strictly applicable to this network; the agreement is empirical rather than analytical. (ii) The bond percolation threshold $p_{c}$ is estimated using the susceptibility peak method (maximum of the mean finite-cluster size), the standard approach in computational percolation [7]. We verify robustness by comparing against the $S=0.5$ crossing and the Molloy–Reed analytical estimate; all three methods confirm $\Delta F_{p}\gg 0$ (Appendix A). (iii) We test sensitivity to threshold heterogeneity using Beta-distributed per-node thresholds centered on $\varphi$ (Appendix D); the compounding vulnerability effect persists under moderate and high heterogeneity, with amplification ranging from $349\%$ to $1{,}955\%$ at $\varphi=0.22$. (iv) The configuration-model approximation in §V.3 ignores BA disassortativity and the correlation structure induced by targeted removal [45]. Appendix C validates the $z_{1}$ framework on configuration-model networks (where it is exact by construction), confirming that the approximation captures the dominant mechanism and that BA-specific correlations partially dampen rather than create the effect. (v) Static networks are studied; adaptive rewiring could mitigate or amplify the effect. (vi) The Watts cascade model is one of several cascade mechanisms; generalization to load-redistribution [18], cascade-based attacks [21], and DebtRank [14] models is an important direction. (vii) The $\kappa>{\sim}10$ threshold is derived for BA networks; the precise threshold depends on both the degree distribution shape and the cascade threshold $\varphi$, and should be understood as a BA-specific onset condition rather than a distribution-independent constant. (viii) In the hub vulnerability experiment (§III), condition B ($\varphi_{\mathrm{hub}}=0.01$) could in principle convert hubs into amplifiers that activate from a single neighbor and immediately pressure k1k-1 others. However, condition Bn ($\varphi_{\mathrm{hub}}=1/k_{i}$, single-neighbor activation without extra amplification) yields statistically indistinguishable cascade sizes (96.5% vs. 95.0%, $p>0.3$), confirming that the effect is firewall removal, not amplification. This equivalence arises because for all hub degrees $k\geq 7$ in our removal set, $\lceil k\cdot 0.01\rceil=\lceil k\cdot(1/k)\rceil=1$—both conditions produce single-neighbor activation. (ix) All cascade simulations use synchronous (parallel) update; asynchronous update may produce quantitatively different cascade sizes, though the qualitative phase structure is expected to persist given the mean-field nature of the $z_{1}$ framework.